OpenStudy (bekkah323):

find an equation of the line having the given slope and containing the given point. m=2/3, (1,-2)

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (bekkah323):

my book told me the answer was y=2/3x+-8/3 but i keep getting y=2/3x+4/3

OpenStudy (anonymous):

take the slope and the point and substitute them into the point slope form then convert that into the slope intercept or standard

OpenStudy (anonymous):

one second while i check

OpenStudy (anonymous):

y+2 = 2/3 ( x-1)

OpenStudy (anonymous):

book is right, double check your steps

OpenStudy (bekkah323):

@completeidiot for the pint slope formula do i only fill in one of the x's and y's because i only have one set coordinates

OpenStudy (anonymous):

yes you fill in y1 and x1 you leave x and y as variables

OpenStudy (bekkah323):

so would it be y--2=2/3(x-1) ?

OpenStudy (anonymous):

yes when you subtract a negative number, it is the same as adding the positive number so your equation will look like y+ 2 = 2/3 (x-1)

OpenStudy (bekkah323):

ok so now what do i do from there?

OpenStudy (anonymous):

you should distribute first

OpenStudy (anonymous):

please write down what you get

OpenStudy (bekkah323):

what do you mean distribute, do you mean the 2/3 to the x and -1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

distribution property of multiplication a(b+c) = ab + ac

OpenStudy (bekkah323):

ok so i got y+2=2/3x -2/3 ( im not very good at adding and subtracting fractions so im not sure if i did that right)

OpenStudy (anonymous):

ok you distributed correctly now you want to convert it to slope intercept form this can be done by subtracting 2 to each side

OpenStudy (anonymous):

however, you cannot directly subtract 2 from -2/3 because they have different denominators this mean you need to convert 2 into a fraction, this can be done by multiplying 1 to it or 3/3

OpenStudy (anonymous):

\[2*1= 2 \] \[1 = \frac{3}{3}\] \[2 * \frac{3}{3} = \frac{ 6}{3}\]

OpenStudy (bekkah323):

so 2/3-2 is 6/3?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[2 = \frac{ 6}{3}\]

OpenStudy (anonymous):

so substituting 2 for 6/3 the equation now looks like y + 6/3 = 2/3 x - 2/3

OpenStudy (anonymous):

now, we subtract 6/3 to both sides

OpenStudy (anonymous):

y = 2/3 x - 2/3 -6/3

OpenStudy (anonymous):

now, \[-\frac{2}{3} - \frac{ 6}{3} = ?\]

OpenStudy (bekkah323):

-8/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

any questions?

OpenStudy (bekkah323):

nope, thank you so much!!!!